OpenStudy (anonymous):

Prove, in complete sentences, whether the left expression is equal to the right expression and discuss which property applies. 6(x − 2) double arrow 6x − 12

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

a(b-c)=ab-ac

OpenStudy (cwrw238):

distributive property

OpenStudy (anonymous):

but are they the same

OpenStudy (cwrw238):

yes they are 6(x - 2) = 6*x + 6*-2 = 6x - 12